henry is 7 times as old as his daughter in five years time he will be 4 times as old as his daughter will be . find their ages
Answer by addingup(3677)

We'll call the father f and the daughter d:
f= 7d We will use this value for f in the next equation:
f+5= 4(d+5)Substitute for f:
7d+5= 4(d+5)Multiply on right to eliminate parenthesis:
7d+5= 4d+20 Subtract 5 on both sides and subtract 4d also on both sides:
3d= 15 Divide both sides by 3:
d= 5 The daughter is 5, and the father is:
f= 7d= 7(5)= 35
----------------------
check:
f+5= 4(d+5)
35+5= 24+16
40= 40 We have the correct answer